# The extinction of the contact process in a one-dimensional random environment with long-range interactions

### v1

@@ Abstract.
We study the contact process on the long-range percolation cluster on $\mathbb{Z}$ where each edge $\langle i,j\rangle$ is open with probability $|i-j|^{-s}$ for $s>2$. Using a renormalization procedure we apply Peierls-type argument to prove that the contact process dies out if the transmission rate is smaller than a critical threshold. Our methods involve the control of crossing probabilities for percolation on randomly-stretched lattices as in [Hilario2023].
@@ 1.1. Model and main results
Let $s>0$. Starting from the complete graph on $\mathbb{Z}$, we declare each unoriented edge $\langle x,y\rangle$ to be open independently with probability $|x-y|^{-s}$, and closed otherwise. The subgraph consisting of only the open edges defines a random graph $G=(\mathbb{Z},\mathcal{E})$, which is almost surely connected since it contains every edge of length one. See Figure 1
@@
Let $\zeta^{A,G}_{t}$ denote the resulting Markov process in $\{0,1\}^{\mathbb{Z}}$, where $1$ stands for infected and 00 for healthy. We may omit the dependence on $G$ and simply write $\zeta^{A}_{t}$. We define the critical infection rate $\lambda_{c}$ as the threshold value for survival. That is,
$\lambda_{c}=\inf\big{\{}\lambda>0\colon\zeta^{\{o\}}_{t}\neq\varnothing\text{ for every }t\geqslant 0\big{\}}.$
Standard ergodicity arguments imply that $\lambda_{c}$ is almost surely constant. Since $G$ almost surely contains a copy of $\mathbb{Z}$, we conclude that $\lambda_{c}<\infty$.
@@
For $G=(\mathbb{Z},\mathcal{E})$, when $s\leqslant 1$, each vertex has infinite degree a.s., therefore the contact processes survives as soon as $\lambda>0$, that is, $\lambda_{c}=0$. In [HaoCan_15], it is shown that if $s$ is sufficiently large, then $\lambda_{c}>0$, and it is conjectured that this should hold as soon as $s>2$. Our main result resolves this conjecture.
@@
During preparation of the final version of this manuscript, we became aware of the recent work [Jahnel25], in which the authors also prove the same conjecture, among other results. Both works were developed independently. Although they share some similarities, the proofs rely on different techniques.
While our results establish the existence of a phase transition for $s>2$, the behavior of the model in the intermediate regime $1<s\leqslant 2$ remains an open question. Determining whether a nontrivial critical infection rate exists in this range would provide a more complete understanding of the model, analogous to the well-established phase diagram for other one-dimensional systems with long-range interactions of order $|i-j|^{-s}$ such as the Ising model [Dyson69, Frolich84] and Bernoulli percolation [Newman86, Aizenman86].
@@ 1.2. Model and definitions
The contact process introduced by Harris in [Harris74] is an interacting particle system often regarded as a model for the spread of an infection (or information) in a graph.
We will study the contact process on $G$ through its graphical representation. We now describe the main ideas involved and refer the reader to the classic work by Harris [Harris78] for a detailed and formal construction.
@@ 1.3. Related works
The contact process on graphs like the $d$-dimensional Euclidean lattice and the $d$-ary regular tree is a well-understood model (see [Liggett1985, Liggett1999] for comprehensive accounts).
@@
Over the past two decades, research has increasingly focused on the contact process on random graphs. To the best of our knowledge, sufficient conditions for the existence of a non-trivial subcritical phase for the contact process on random graphs with unbounded degree distributions were first established by Ménard and Singh [Menard_16]. Their conditions apply to graphs such as Delaunay triangulations and random geometric graphs. In [HaoCan_15], Can verified these conditions for the long-range percolation graph when $s>102$. More recently, Jahnel et al. [Jahnel25] established the existence of a subcritical phase for all $s>2$, independently of our work. Their proof relies on a coupling with a random walk on a random environment while ours, as is classic in percolation, is based on a Peierls argument.
@@
In [Jahnel25], the authors also study the phase transition for the contact process on one-dimensional random geometric graphs with i.i.d. random radii. They show that a phase transition takes place if, and only if, the radius distribution has finite mean. Our methods work for that kind of graphs as long as the radius has tail distribution decaying as $r^{1+\alpha}$ for any $\alpha>0$. The contact process on the dynamic version of this random graph, where the radii update independently at rate one was studied in [GL2022].
@@
Many other authors have considered different kinds of random graphs, mainly Galton-Watson trees and the configuration model. We now list some of the results in the field and refer the reader to [Valesin2024] for a gentle and comprehensive account on the topic.
@@
In [Pemantle01], Pemantle and Stacey studied the contact process on Galton-Watson trees with bounded-degree offspring distributions. They identified examples where the critical points for local and global survival differ, leading to a double phase transition, as observed in regular $d$-ary trees. However, the existence of a subcritical phase for Galton-Watson trees with unbounded-degree offspring remained an open question. Huang and Durrett [Huang2020] showed that if the offspring distribution is subexponential, no subcritical phase exists. In [Bhamidi2021], Bhamidi et al. established that a subcritical phase exists if and only if the offspring distribution has exponential tails.
@@
The contact process on random graphs generated by the configuration model on $n$ vertices with a power-law degree distribution (with exponent greater than 3) was studied by Chatterjee and Durrett [ChatterjeeDurrett2009]. As these graphs are finite, the contact process eventually dies out. However, they proved that no phase transition occurs, meaning that for any positive infection parameter $\lambda$, the process survives for an exceptionally long time. Specifically, the extinction time grows at least as a stretched exponential function of the number of vertices. Mountford et al. [Mountford2016] improved this result, showing that the survival time is exponential in the number of vertices. Additionally, Mountford et al. [MountfordValesinYao2013] established sharp bounds for the density of infected sites. For the particular case when the degrees equal some constant $d$, also called random $d$-regular graph, Mourrat and Valesin [Mourrat2016] and Lalley and Su [Lalley2017] demonstrated the existence of a phase transition. The latter work also established a cutoff for the density of infected sites. Bhamidi et al. [Bhamidi2021] showed that a phase transition occurs if and only if the degree distribution has an exponential moment.
@@ 1.4. Summary of the paper
This is the content of Section 3.
Peierls argument: A Peierls-type argument shows that the contact process dies out if there exists a semicircuit of good boxes surrounding the origin. In fact, since the left and right boundaries of each box project onto cut-points, if the infection ever enters a good box $V_{k}\times[jT,(j+1)T)$, it cannot propagate to neighboring boxes $V_{k-1}\times[jT,(j+1)T)$, $V_{k+1}\times[jT,(j+1)T)$, or $V_{k}\times[(j+1)T,(j+2)T)$.
@@
Comparison with bond percolation on a randomly stretched lattice: It turns out that the probability of finding semicircuits in the above site percolation model can be compared with that of a bond percolation model studied in [Hilario2023].
@@
(1) $p_{e}=\left\{\begin{array}[]{ll}p,&\text{if }e=\{(i,j),(i,j+1)\}\text{ for some }i,j,\\ p^{\xi_{i+1}},&\text{if }e=\{(i,j),(i+1,j)\}\text{ for some }i,j.\end{array}\right.$
When the sequence $(\xi_{n})_{n}$ consists of independent geometric random variables, the model is reminiscent of the one introduced in [JMP] and studied further in [Hoffman_2005, Hilario2023+]. In our setting, however, the marginal distribution of $\xi_{n}$ is no longer geometric, requiring a different analytical approach. To handle this, we rely on the results developed in [Hilario2023]. There the authors derive estimates for the probability of crossing rectangles and apply them to show that for $p$ large enough the connected component containing the origin is infinite with positive probability. Here we use crossings of rectangles in order to build a semicircuit around the origin with high probability. This is done in Sections 4 and 5.
@@ Acknowledgements
The authors thank Daniel Valesin for valuable discussions, comments and suggestions. P.G. acknowledges the hospitality of EPFL, where discussions with Prof. Thomas Mountford were initially held. The research of P.G. is partially supported by FAPESP grants 21/10170-7 and 23/13453-5. The research of M.H. is partially supported by FAPEMIG grant APQ-01214-21 (Universal), CNPq grant 312566/2023-9 (Produtividade em Pesquisa) and CNPq grant 406001/2021-9 (Universal). B.N.B.L. is partially supported by CNPq grant 315861/2023-1.
@@ 2. Control of the environment, cut-points
[…]
@@
(5) (o is a $\begin{split}\mathbb{P}(\text{$o$ is a cut-point})&=\prod_{i\geqslant 1}\prod_{j\geqslant 1}\mathbb{P}\big{(}\langle-i,j\rangle\textrm{ is closed}\big{)}\\ &=\prod_{i\geqslant 1}\prod_{j\geqslant 1}\left(1-\frac{1}{|j+i|^{s}}\right)>0,\end{split}$
@@
Given the above construction of the sequence $\{x_{k},\leavevmode\nobreak\ k\in\mathbb{Z}\setminus\{0\}\}$, one can readily check that $\{\theta_{k}\colon k\in\mathbb{Z}\setminus\{0\}\}$ forms an i.i.d. sequence. We denote by $\theta$ a random variable with the same distribution as any of the variables in that sequence. The following result shows that $\theta$ has some moment of order larger than one.
@@
For every $20\leqslant\delta<s-$ 2,
@@
(It is not difficult to verify that the $min$ and $max$ run over nonempty sets and are well-defined.)
@@
This gives rise to a sequence $\{v_{k},\leavevmode\nobreak\ k\in\mathbb{Z}\setminus\{0\}\}$ of strong cut-points spanning $\mathbb{Z}$.
@@
Given the new sequence of strong cut-points $\{v_{k},\leavevmode\nobreak\ k\in\mathbb{Z}\setminus\{0\}\}$, we will partition $\mathbb{Z}$ into intervals $V_{k}$, where
@@
$\mathcal{E}_{k}=\begin{cases}\bigl{\{}\langle x,y\rangle\in\mathcal{E}\colon\max\{x,y\}\in V_{k}\bigr{\}},&\leavevmode\nobreak\ \textrm{ if }k\in{\mathbb{Z}}_{-},\\ \big{\{}\langle x,y\rangle\in\mathcal{E}\colon\min\{x,y\}\in V_{k}\big{\}},&\leavevmode\nobreak\ \textrm{ if }k\in{\mathbb{Z}}_{+}.\end{cases}$
When $k$ is odd, $|\mathcal{E}_{k}|=1$. In fact, since $v_{k}$ is a strong cut-point, either $\mathcal{E}_{2k-1}=\{\langle v_{k},v_{k}+1\rangle\}$ or $\mathcal{E}_{2k+1}=\{\langle v_{k}-1,v_{k}\rangle\}$ according to whether $k$ is positive or negative. The random variables $\{|\mathcal{E}_{k}|,\leavevmode\nobreak\ k\in 2\mathbb{Z}\setminus\{0\}\}$ are i.i.d. The next result shows that they have finite moments of order greater than one.
@@
$x_{1}^{L}=\min\{x>x_{+}\colon\textrm{ $\langle i,j\rangle$ is closed for all $0\leqslant i<x<j$ with $j-i>L$}\},$
@@
$x_{n+1}^{L}=\min\{x>x_{n}^{L}\colon\textrm{ $\langle i,j\rangle$ is closed forall $x_{n}^{L}\leqslant i<x<j$ with $j-i>L$}\}.$
Similarly, for $x_{n}^{L}$, with $n\leqslant-1$. We recover the original sequence $(x_{k})_{k\in\mathbb{Z}\setminus\{0\}}$ simply by taking $L=1$. Writing,
@@
Let $20<\delta<s-$ 2. There exists $L_{o}=L_{o}(s,\delta)$ such that, for every $L\geqslant L_{o}(s,\delta)$,
@@
Before we proceed to prove Lemma 1 let us introduce the backward exploration, reminiscent of a procedure found in [GGNR, Section 2].
@@
For the sum appearing in the right-hand side, ignoring $k=1$ and $k=n-1$ we obtain the upper bound
@@
[…]
@@
Including the terms $k=1$ and $k=n-1$ which are equal to $1/(n-1)^{s-1}$, we get
@@
On the event $\{\theta_{1}^{L}\geqslant n,\leavevmode\nobreak\ x^{L}_{1}=z\}$ we have that, for every $z^{\prime}\in(z,z+n]$,
@@
otherwise we would obtain $x_{k+1}^{L}<z^{\prime}\leqslant z+n$, hence $\theta_{k}^{L}<n$. Therefore, on the same event, we can define recursively a sequence by setting $z_{1}=z+n$, and for $\ell\geqslant 1$, we define
@@
$\begin{split}\mathbb{P}\left(\theta_{1}^{L}\geqslant n\leavevmode\nobreak\ |\leavevmode\nobreak\ x_{1}^{L}=z\right)&\leqslant\mathbb{P}\left(\exists\leavevmode\nobreak\ m\geqslant 1\textrm{ such that }z_{m+1}=z\right)\\ &=\mathbb{P}\Big{(}\exists\leavevmode\nobreak\ m\geqslant 1\colon\sum_{k=1}^{m}Y_{k}=-n\Big{)}.\end{split}$
@@
(16) $\begin{split}\mathbb{P}\Big{(}\exists\leavevmode\nobreak\ m\geqslant 1\colon\sum_{k=1}^{m}Y_{k}=-n\Big{)}&\leqslant\sum_{m=1}^{M}\mathbb{P}\Big{(}\sum_{k=1}^{m}Y_{k}=-n\Big{)}+\sum_{m>M}\mathbb{P}\Big{(}\sum_{k=1}^{m}Y_{k}=-n\Big{)}\\ &<\sum_{m=1}^{M}\frac{c_{o}^{m}}{n^{s-1}}+\sum_{m>M}(1-a)^{m}\\ &<\frac{c_{o}^{M}}{n^{s-1}}+\frac{(1-a)^{M}}{a}<\frac{c}{n^{1+\delta}}.\end{split}$
@@
Write $A=\{x_{k}\colon k\in\mathbb{Z}\setminus\{0\}\}$ and $A_{L}=\{x_{k}^{L}\colon k\in\mathbb{Z}\setminus\{0\}\}$. Because $A\subset A_{L}$, for every $k\in{\mathbb{Z}}^{+}$ and $z\in\mathbb{Z}_{+}$, we can write
@@
$\kappa=\mathbb{P}\big{(}\text{$\langle-i,j\rangle$ is closed for every $i,j>0$such that $i+j\leqslant L$}\big{)}>0.$
@@
is independent of $\{X_{i}^{L},i\leqslant m\}$ and is distributed as a geometric random variable with parameter $\kappa$. Writing $m^{\prime}=m+NL$ and $m^{\ast}=\min\{i>m\colon x_{i}^{L}\in A\}$, then $m^{\ast}\leqslant m^{\prime}$. Therefore, defining $\zeta=x_{m^{\prime}}^{L}-x_{m}^{L}$, on the event $\{x_{m}^{L}=x_{k}=z\}$, we have
@@
Therefore, $\mathbb{P}(\theta_{k}\geqslant n\leavevmode\nobreak\ |\leavevmode\nobreak\ x_{m}^{L}=x_{k}=z)\leqslant\mathbb{P}(\zeta\geqslant n\leavevmode\nobreak\ |\leavevmode\nobreak\ x_{m}^{L}=x_{k}=z)$.
@@
$\zeta=\sum_{i=m}^{m+LN-1}X_{i}^{L}.$
Now, if $\{W_{i},\leavevmode\nobreak\ i\geqslant 1\}$, are i.i.d. random variables with the same distribution as $X_{i}^{L}$, and $G$ is a geometric random variable with parameter $\kappa$ independent of all the rest, then the conditional distribution of $\zeta$ given $\{x_{m}^{L}=x_{k}=z\}$ is the same as
@@
$\mathbb{P}\left(\theta_{k}\geqslant n\leavevmode\nobreak\ |\leavevmode\nobreak\ x_{m}^{L}=x_{k}=z\right)\leqslant\mathbb{P}\left(\zeta\geqslant n\leavevmode\nobreak\ |\leavevmode\nobreak\ x_{m}^{L}=x_{k}=z\right)=\mathbb{P}\left(W\geqslant n\right).$
@@
$\mathbb{P}(\theta\geqslant n)$ $=\mathbb{P}\left(\theta_{k}\geqslant n\leavevmode\nobreak\ |\leavevmode\nobreak\ x_{k}=z\right)$
$=\frac{1}{\mathbb{P}(x_{k}=z)}\sum_{m\geqslant k}\left[\mathbb{P}\left(\theta_{k}\geqslant n\leavevmode\nobreak\ |\leavevmode\nobreak\ x_{m}^{L}=x_{k}=z\right)\mathbb{P}\left(x_{m}^{L}=x_{k}=z\right)\right]$
@@
$\mathbb{P}(\tilde{\theta}=n)=\mathbb{P}(\theta_{1}+\theta_{2}=n\leavevmode\nobreak\ |\leavevmode\nobreak\ \theta_{1}+\theta_{2}\neq 2),\leavevmode\nobreak\ n\geqslant 3.$

### v2

@@ Abstract.
We study the contact process on the long-range percolation cluster on $\mathbb{Z}$ where each edge $\langle i,j\rangle$ is open with probability $|i-j|^{-s}$ for $s>2$. Using a renormalization procedure we apply Peierls-type argument to prove that the contact process dies out if the transmission rate is smaller than a critical threshold. Our methods involve the control of crossing probabilities for percolation on randomly-stretched lattices as in [14].
@@ 1.1. Model and main results
Let $s>0$. Starting from the complete graph on $\mathbb{Z}$, we declare each unoriented edge $\langle x,y\rangle$ to be open independently with probability $|x-y|^{-s}$, and closed otherwise. The subgraph consisting of only the open edges defines a random graph $G=(\mathbb{Z},\mathcal{E})$, which is almost surely connected since it contains every edge of length one. See Figure 1.
@@
Let $\zeta^{A,G}_{t}$ denote the resulting Markov process in $\{0,1\}^{\mathbb{Z}}$, where $1$ stands for infected and 0 for healthy. We may omit the dependence on $G$ and simply write $\zeta^{A}_{t}$. We define the critical infection rate $\lambda_{c}$ as the threshold value for survival. That is,
$\lambda_{c}(\mathcal{E})=\inf\big\{\lambda>0\colon{P}_{\lambda}^{\mathcal{E}}\big(\zeta^{\{o\}}_{t}\neq\varnothing\text{ for every }t\geqslant 0\big)>0\big\},$
where $P^{\mathcal{E}}_{\lambda}$ stands for the law of the contact process on $\mathcal{E}$. Standard ergodicity arguments imply that $\lambda_{c}(\mathcal{E})$ does not depend on the realization of $\mathcal{E}$, almost surely; hence we write simply $\lambda_{c}$. Since $G$ almost surely contains a copy of $\mathbb{Z}$, we conclude that $\lambda_{c}<\infty$; this was proved originally by [12] (see Section 1 of Chapter VI of [20] or Chapter 6 of [10] for a more comprehensive approach).
@@
For $G=(\mathbb{Z},\mathcal{E})$, when $s\leqslant 1$, each vertex has infinite degree a.s., therefore the contact processes survives as soon as $\lambda>0$, that is, $\lambda_{c}=0$. In [4], it is shown that if $s$ is sufficiently large, then $\lambda_{c}>0$, and it is conjectured that this should hold as soon as $s>2$. Our main result resolves this conjecture.
@@
During preparation of the final version of this manuscript, we became aware of the recent work [17], in which the authors also prove the same conjecture, among other results. Both works were developed independently. Although they share some similarities, the proofs rely on different techniques.
While our results establish the existence of a phase transition for $s>2$, the behavior of the model in the intermediate regime $1<s\leqslant 2$ remains an open question. Determining whether a nontrivial critical infection rate exists in this range would provide a more complete understanding of the model, analogous to the well-established phase diagram for other one-dimensional systems with long-range interactions of order $|i-j|^{-s}$ such as the Ising model [6, 7] and Bernoulli percolation [26, 1].
@@ 1.2. Model and definitions
The contact process introduced by Harris in [12] is an interacting particle system often regarded as a model for the spread of an infection (or information) in a graph.
We will study the contact process on $G$ through its graphical representation. We now describe the main ideas involved and refer the reader to the classic work by Harris [13] for a detailed and formal construction.
@@ 1.3. Related works
The contact process on graphs like the $d$-dimensional Euclidean lattice and the $d$-ary regular tree is a well-understood model (see [20, 21] for comprehensive accounts).
@@
Over the past two decades, research has increasingly focused on the contact process on random graphs. To the best of our knowledge, sufficient conditions for the existence of a non-trivial subcritical phase for the contact process on random graphs with unbounded degree distributions were first established by Ménard and Singh [22]. Their conditions apply to graphs such as Delaunay triangulations and random geometric graphs. In [4], Can verified these conditions for the long-range percolation graph when $s>102$. More recently, Jahnel et al. [17] established the existence of a subcritical phase for all $s>2$, independently of our work. Their proof relies on a coupling with a random walk on a random environment while ours, as is classic in percolation, is based on a Peierls argument.
@@
In [17], the authors also study the phase transition for the contact process on one-dimensional random geometric graphs with i.i.d. random radii. They show that a phase transition takes place if, and only if, the radius distribution has finite mean. The contact process on the dynamic version of this random graph, where the radii update independently at rate one was studied in [9].
@@
Many other authors have considered different kinds of random graphs, mainly Galton-Watson trees and the configuration model. We now list some of the results in the field and refer the reader to [28] for a gentle and comprehensive account on the topic.
@@
In [27], Pemantle and Stacey studied the contact process on Galton-Watson trees with bounded-degree offspring distributions. They identified examples where the critical points for local and global survival differ, leading to a double phase transition, as observed in regular $d$-ary trees. However, the existence of a subcritical phase for Galton-Watson trees with unbounded-degree offspring remained an open question. Huang and Durrett [5] showed that if the offspring distribution is subexponential, no subcritical phase exists. In [2], Bhamidi et al. established that a subcritical phase exists if and only if the offspring distribution has exponential tails.
@@
The contact process on random graphs generated by the configuration model on $n$ vertices with a power-law degree distribution (with exponent greater than 3) was studied by Chatterjee and Durrett [3]. As these graphs are finite, the contact process eventually dies out. However, they proved that no phase transition occurs, meaning that for any positive infection parameter $\lambda$, the process survives for an exceptionally long time. Specifically, the extinction time grows at least as a stretched exponential function of the number of vertices. Mountford et al. [24] improved this result, showing that the survival time is exponential in the number of vertices. Additionally, Mountford et al. [23] established sharp bounds for the density of infected sites. For the particular case when the degrees equal some constant $d$, also called random $d$-regular graph, Mourrat and Valesin [25] and Lalley and Su [19] demonstrated the existence of a phase transition. The latter work also established a cutoff for the density of infected sites. Bhamidi et al. [2] showed that a phase transition occurs if and only if the degree distribution has an exponential moment.
@@ 1.4. Summary of the paper
This is the content of Section 3. The reader is invited to consult Figure 4 for an illustration of a good box.
Peierls argument: A Peierls-type argument shows that the contact process dies out if there exists a semicircuit of good boxes surrounding the origin. In fact, since the left and right boundaries of each box project onto cut-points, if the infection ever enters a good box $V_{k}\times[jT,(j+1)T)$, it could only propagate to neighboring boxes $V_{k-1}\times[jT,(j+1)T)$, $V_{k+1}\times[jT,(j+1)T)$, or $V_{k}\times[(j+1)T,(j+2)T)$ passing through their common boundaries. However that propagation if forbidden if the neighboring boxes are also good (see Figure 4). Therefore, an infection starting at the origin cannot reach the exterior of the semicircuit.
@@
Comparison with bond percolation on a randomly stretched lattice: It turns out that the probability of finding semicircuits in the above site percolation model can be compared with that of a bond percolation model studied in [14].
@@
[…]
We will write $\mathbb{P}^{\Lambda}_{\rho}$ for the corresponding probability measure induced in $\{0,1\}^{\mathcal{E}(\mathbb{Z}^{d})}$
When the sequence $(\xi_{n})_{n}$ consists of independent geometric random variables, the model is reminiscent of the one introduced in [18] and studied further in [16, 15]. In our setting, however, the marginal distribution of $\xi_{n}$ is no longer geometric, requiring a different analytical approach. To handle this, we rely on the results developed in [14]. There the authors derive estimates for the probability of crossing rectangles and apply them to show that for $p$ large enough the connected component containing the origin is infinite with positive probability. Here we use crossings of rectangles in order to build a semicircuit around the origin with high probability. This is done in Sections 4 and 5.
@@ Acknowledgements
The authors thank Daniel Valesin for valuable discussions, and two referees for comments, suggestions and corrections. P.G. acknowledges the hospitality of EPFL, where discussions with Prof. Thomas Mountford were initially held. The research of P.G. is partially supported by Fundação de Amparo à Pesquisa do Estado de São Paulo grants 21/10170-7 and 23/13453-5. The research of M.H. is partially supported by Fundação de Amparo à Pesquisa do Estado de Minas Gerais grant APQ-01214-21 (Universal), Conselho Nacional de Desenvolvimento Científico e Tecnológico grant 312566/2023-9 (Produtividade em Pesquisa) and grant 406001/2021-9 (Universal). B.N.B.L. is partially supported by Conselho Nacional de Desenvolvimento Científico e Tecnológico grant 315861/2023-1.
@@ 2. Control of the environment, cut-points
(4) $x_{-}$ $=$ $\inf\big\{i\in{\mathbb{Z}}_{-}\colon\text{ $\langle i,j\rangle$ is open for some $j\in{\mathbb{Z}}_{+}$}\big\},$
@@
(5) $\begin{split}\mathbb{P}(\text{$o$ is a cut-point})&=\prod_{i\geqslant 1}\prod_{j\geqslant 1}\mathbb{P}\big(\langle-i,j\rangle\textrm{ is closed}\big)=\prod_{i\geqslant 1}\prod_{j\geqslant 1}\left(1-\frac{1}{|j+i|^{s}}\right)\\ &\geqslant\tfrac{1}{2}\exp\left(-\sum_{i\geqslant 1}\sum_{j\geqslant 1}\frac{1}{|j+i|^{s}}\right)>0,\end{split}$
@@
Given the above construction of the sequence $\{x_{k},\penalty 10000\ k\in\mathbb{Z}\setminus\{0\}\}$, one can readily check that $\{\theta_{k}\colon k\in\mathbb{Z}\setminus\{0\}\}$ forms an i.i.d. sequence. We denote by $\theta$ a random variable with the same distribution as any of the variables in that sequence. The following result shows that $\theta$ has some moment of order larger than one.
@@
For every $0\leqslant\delta<s-2$,
@@
(It is not difficult to verify that the minmin and maxmax run over nonempty sets and are well-defined.)
@@
This gives rise to a sequence $\{v_{k},\penalty 10000\ k\in\mathbb{Z}\setminus\{0\}\}$ of strong cut-points spanning $\mathbb{Z}$.
@@
Given the new sequence of strong cut-points $\{v_{k},\penalty 10000\ k\in\mathbb{Z}\setminus\{0\}\}$, we will partition $\mathbb{Z}$ into intervals $V_{k}$, where
@@
See Figure 3 for an illustration of these sets.
$v_{k-1}$$v_{k}$$v_{k+1}$$v_{k+2}$$V_{2k-2}$$V_{2k}$$V_{2k+2}$
Figure 3. The representation of the strong cut-points $v_{k}$ and the corresponding sets $V_{2k}$.
@@
[…]
When $k$ is odd, $|\mathcal{E}_{k}|=1$. In fact, since $v_{k}$ is a strong cut-point, either $\mathcal{E}_{2k-1}=\{\langle v_{k},v_{k}+1\rangle\}$ or $\mathcal{E}_{2k+1}=\{\langle v_{k}-1,v_{k}\rangle\}$ according to whether $k$ is positive or negative. The random variables $\{|\mathcal{E}_{k}|,\penalty 10000\ k\in 2\mathbb{Z}\setminus\{0\}\}$ are i.i.d. The next result shows that they have finite moments of order greater than one.
@@
[…]
@@
[…]
Similarly, for $x_{n}^{L}$, with $n\leqslant-1$.
We recover the original sequence $(x_{k})_{k\in\mathbb{Z}\setminus\{0\}}$ simply by taking $L=1$. Writing,
@@
Let $0<\delta<s-2$. There exists $L_{o}=L_{o}(s,\delta)$ such that, for every $L\geqslant L_{o}(s,\delta)$,
@@
Before we proceed to prove Lemma 1 let us introduce the backward exploration, reminiscent of a procedure found in [8, Section 2].
@@
For the sum appearing in the right-hand side, ignoring $k=1$ and k=n1k=n-1 we obtain the upper bound
@@
$<\frac{2}{n^{2s-3}}\int_{\frac{1}{n}}^{\frac{1}{2}}\frac{1}{x^{s-1}(1-x)^{s-1}}dx$
$<\frac{2}{n^{2s-3}}\int_{\frac{1}{n}}^{\frac{1}{2}}\frac{1}{x^{s-1}(1/2)^{s-1}}dx$
@@
Including the terms $k=1$ and k=n1k=n-1 which are equal to $1/(n-1)^{s-1}$, we get
@@
On the event $\{\theta_{1}^{L}\geqslant n,\penalty 10000\ x^{L}_{1}=z\}$ we have that, for every $z^{\prime}\in(z,z+n]$,
@@
otherwise we would obtain $x_{2}^{L}<z^{\prime}\leqslant z+n$, hence $\theta_{1}^{L}<n$. Therefore, on the same event, we can define recursively a sequence by setting $z_{1}=z+n$, and for $\ell\geqslant 1$, we define
@@
$\begin{split}\mathbb{P}\left(\theta_{1}^{L}\geqslant n\penalty 10000\ |\penalty 10000\ x_{1}^{L}=z\right)&\leqslant\mathbb{P}\left(\exists\penalty 10000\ m\geqslant 1\textrm{ such that }z_{m+1}=z\right)\\ &=\mathbb{P}\Big(\exists\penalty 10000\ m\geqslant 1\colon\sum_{k=1}^{m}Y_{k}=-n\Big).\end{split}$
@@
(16) $\begin{split}\mathbb{P}\Big(\exists\penalty 10000\ m\geqslant 1\colon\sum_{k=1}^{m}Y_{k}=-n\Big)&\leqslant\sum_{m=1}^{M}\mathbb{P}\Big(\sum_{k=1}^{m}Y_{k}=-n\Big)+\sum_{m>M}\mathbb{P}\Big(\sum_{k=1}^{m}Y_{k}=-n\Big)\\ &<\sum_{m=1}^{M}\frac{c_{o}^{m}}{n^{s-1}}+\sum_{m>M}(1-a)^{m}\\ &<\frac{c_{o}^{M}}{n^{s-1}}+\frac{(1-a)^{M}}{a}<\frac{c}{n^{1+\delta}}.\end{split}$
@@
Write $A=\{x_{k}\colon k\in\mathbb{Z}\setminus\{0\}\}$ and $A_{L}=\{x_{k}^{L}\colon k\in\mathbb{Z}\setminus\{0\}\}$. Because $A\subset A_{L}$, for every $k\in\mathbb{Z}_{+}$ and $z\in\mathbb{Z}_{+}$, we can write
@@
[…]
@@
is independent of $\{\theta_{i}^{L},i\leqslant m\}$ and is distributed as a geometric random variable with parameter $\kappa$. Writing $m^{\prime}=m+NL$ and $m^{\ast}=\min\{i>m\colon x_{i}^{L}\in A\}$, then $m^{\ast}\leqslant m^{\prime}$. Therefore, defining $\zeta=x_{m^{\prime}}^{L}-x_{m}^{L}$, on the event $\{x_{m}^{L}=x_{k}=z\}$, we have
@@
Therefore, $\mathbb{P}(\theta_{k}\geqslant n\penalty 10000\ |\penalty 10000\ x_{m}^{L}=x_{k}=z)\leqslant\mathbb{P}(\zeta\geqslant n\penalty 10000\ |\penalty 10000\ x_{m}^{L}=x_{k}=z)$.
@@
$\zeta=\sum_{i=m}^{m+LN-1}\theta_{i}^{L}.$
Now, if $\{W_{i},\penalty 10000\ i\geqslant 1\}$, are i.i.d. random variables with the same distribution as $\theta_{i}^{L}$, and $G$ is a geometric random variable with parameter $\kappa$ independent of all the rest, then the conditional distribution of $\zeta$ given $\{x_{m}^{L}=x_{k}=z\}$ is the same as
@@
$\mathbb{P}\left(\theta_{k}\geqslant n\penalty 10000\ |\penalty 10000\ x_{m}^{L}=x_{k}=z\right)\leqslant\mathbb{P}\left(\zeta\geqslant n\penalty 10000\ |\penalty 10000\ x_{m}^{L}=x_{k}=z\right)=\mathbb{P}\left(W\geqslant n\right).$
@@
$\mathbb{P}(\theta_{k}\geqslant n)$ $=\mathbb{P}\left(\theta_{k}\geqslant n\penalty 10000\ |\penalty 10000\ x_{k}=z\right)$
$=\frac{1}{\mathbb{P}(x_{k}=z)}\sum_{m\geqslant k}\left[\mathbb{P}\left(\theta_{k}\geqslant n\penalty 10000\ |\penalty 10000\ x_{m}^{L}=x_{k}=z\right)\mathbb{P}\left(x_{m}^{L}=x_{k}=z\right)\right]$
@@
$\mathbb{P}(\tilde{\theta}=n)=\mathbb{P}(\theta_{1}+\theta_{2}=n\penalty 10000\ |\penalty 10000\ \theta_{1}+\theta_{2}\neq 2),\penalty 10000\ n\geqslant 3.$